Firm D has net income of $61,750, sales of $1,625,000 and average total assets of $650,000. Calculate the firm’s margin, turnover and ROI
Answer #1

margin= net income/Sale

=(61750/1,625,000)=3.8%

Turnover=Sales/Average total assets

=(1625000/650000)=2.5

ROI=Net income/Average total assets

=61750/650,000=9.5%